OpenStudy (anonymous):

The sum of 12 times the larger number and 11 times the smaller is -36. The difference of 12 times the larger and 7 times the smaller is 36. Find the numbers.

OpenStudy (helder_edwin):

let \(x\) and \(y\) the two numbers we r looking for. and suppose that \(x>y\). then \[ \large 12x+11y=-36 \] and \[ \large 12x-7y=36 \]

OpenStudy (anonymous):

alright then should I move forward with the addition method?

OpenStudy (helder_edwin):

u have a linear system \[ \large \begin{cases} 12x+11y=-36\\ 12x-7y=36 \end{cases} \]

OpenStudy (helder_edwin):

use whichever method u want to solve it.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

thanks

OpenStudy (helder_edwin):

u r welcome